OpenStudy (anonymous):

Solve sin^-1 X - cos^-1 X = sin^-1 (3x-2)?

OpenStudy (anonymous):

@satellite73 @lgbasallote @hartnn @Shane_B @Mimi_x3

OpenStudy (mimi_x3):

please write it properly \[ sin^{-1}x - cos^{-1}x = sin^{-1}{3x-2}\]

OpenStudy (mimi_x3):

Is that right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

@Mimi_x3 i dont knw LATEX......

OpenStudy (anonymous):

this one is much easier

OpenStudy (anonymous):

\(\sin^{-1}(x)+\cos^{-1}(x)=\frac{\pi}{2}\) so \[\cos^{-1}(x)=\frac{\pi}{2}-\sin^{-1}(x)\] and you have \[2\sin^{-1}(x)-\frac{\pi}{2}=\sin^{-1}(3x-2)\] hmm now i am stuck

OpenStudy (anonymous):

for some reason, this is the same as \[\cos^{-1}(2-3x)=2\sin^{-1}(x)\] but i am not sure why this i can solve

OpenStudy (anonymous):

ok i got that it is because \[\sin^{-1}(3x-2)+\cos^{-1}(3x-2)=\frac{\pi}{2}\] and so \[\sin^{-1}(3x-2)=\frac{\pi}{2}-\cos(3x-2)=\frac{\pi}{2}+\cos^{-1}(2-3x)\]

OpenStudy (anonymous):

so on the right we get \[\frac{\pi}{2}+\cos^{-1}(2-3x)\] on the left we can write \[\sin^{-1}(x)-\cos^{-1}(x)=2\sin(x)+\frac{\pi}{2}\] since \[\cos^{-1}(x)=\frac{\pi}{2}-\sin^{-1}(x)\] damn i messed up somewhere

OpenStudy (anonymous):

well, if you can figure out why this says \[2\sin^{-1}(x)=\cos^{-1}(2-3x)\] then it is easy to solve take the cosine of both sides and you will get \[1-2x^2=2-3x\] solve and get \(x=1\) or \(x=\frac{1}{2}\)

OpenStudy (anonymous):

Yupp....@satellite73 u r correct